How do I find the maximum height of a baseball which is hit with an upward velocity of 90 feet per second when the initial height of the ball was 3 feet?
Answer by ankor@dixie-net.com(22740)

How do I find the maximum height of a baseball which is hit with an upward velocity of 90 feet per second when the initial height of the ball was 3 feet?
:
Use the equation: height = -16t^2 + 90t + 3; where t is the time in seconds
:
Use the vertex formula x = -b/(2a)
:
In our equation a = -16 and b = 90
:
t = -90)/2(-16)
t = -90/-32
t = +2.8125 sec when the ball reaches max height
:
Substitute 2.8125 for t in the original equation to find the height:
:
h = -16(2.8125^2) + 90(2.8125) + 3
h = -16(7.91) + 253.125 + 3
h = -126.56 + 253.125 + 3
h = 129.565 ft is the max height
: